4. A student stays at home.  Let event N = the student watches Netflix. Let event Y = the student watches the very educational youtube videos made by her/his instructor. Suppose P(N) = 0.1, P(Y) = 0.8, and P(N|Y) = 0.2. Find P(N and Y).

A. 0.1 B. 0.16 C. 0.18 D. 0.21 E.0.28
Answer #1

since P(N|Y) =P(N and Y) /P(Y)

therefore from given details : P(N and Y) =P(Y)*P(N|Y) =0.8*0.2 =0.16
option b is correct